OpenStudy (anonymous):

One maid can clean the house in 4hr. Another maid can clean in 2hr. How long will it take if they do the job together?

OpenStudy (anonymous):

\[\frac{2\times 4}{2+4}\] \[\frac{8}{6}\] \[\frac{4}{3}\]

OpenStudy (anonymous):

one hour and a third one hour and twenty minutes